OpenStudy (anonymous):

use log 5~0.699 and log 12~1.079 to approximate the value of log 25

Parth (parthkohli):

You don't really need \(\log(12)\).\[\log(25)\]\[= \log(5\times 5)\]\[= \log(5) + \log(5)\]\[= 0.699 + 0.699\]\[= 1.398\]

OpenStudy (anonymous):

how would u condense the expression 3 log x + log 7

OpenStudy (anonymous):

\[\log \left( 5 \right) + \log \left( 5 \right)\] \[1.609 + 1.609 \approx 3.218\]

Parth (parthkohli):

@some_someone Err, no, that is the natural logarithm. Here we're talking about the base-10 which is mentioned in the question.

OpenStudy (anonymous):

oh okaii it wasnt clear..

OpenStudy (anonymous):

how would u condense the expression 3 log x + log 7